# Bank–Laine functions with preassigned number of zeros

### v2

@@ 4.2. Definition of a quasiregular map
we will consider the functions
(4.19) $\bar{g}_{m_{k},n_{k}}(z)=g_{m_{k},n_{k}}(z)$
@@
(4.20) $\bar{g}_{m_{k},n_{k}}(z)=\frac{g_{m_{k},n_{k}}(z)+1}{2}$
@@
$\check{\Pi}_{k}=\{z=x+iy:\quad 2\pi(-k)\leq y\leq 2\pi(-k+1)\}$
we will consider the functions
(4.21) $\bar{g}_{m_{k},n_{k}}(z)=g_{m_{k},n_{k}}(z).$
In view of the explicit expressions of $\bar{g}_{m_{k},n_{k}}$ defined in (4.19), (4.20) and (4.21), the maps $U$, $V$ and $G$ will not commute with complex conjugation. However, we shall only define them in the upper half-plane since the definitions in the lower half-plane are similar.
@@
$\begin{split}\Pi_{k}^{+}=\{x+iy:x>0,\quad 2\pi(k-1)<y<2\pi k\}.\end{split}$
@@ 4.3. Completion of the proof
With the definition of $G$ in (4.22) in previous section, here we first prove Theorem 1.6 under the assumption that $\rho\in(1/2,1)$. Now, we can follow exactly the same process as that in [5, Section 3.3] to estimate the complex dilation of $G$ to prove that $G$ is indeed quasiregular in the plane and also that $G$ satisfies the hypothesis of the Teichmüller–Wittich–Belinskii theorem. This theorem, together with the existence theorem for quasiconformal mappings, yields that there exists a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ and a meromorphic function $F$ such that
@@
Then $F=G\circ\tau^{-1}$ is a locally univalent meromorphic function. Define $E=F/F^{\prime}$. Then $E$ is a Bank–Laine function. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A(z)$.
We begin to estimate the numbers of poles and zeros of $F$ respectively. Again, two positive quantities $S_{1}(r)$ and $S_{2}(r)$ are written as $S_{1}(r)\asymp S_{2}(r)$ if there are two positive constants $d_{1}$ and $d_{2}$ such that $d_{2}S_{2}(r)\leq S_{1}(r)\leq d_{1}S_{2}(r)$ for all large $r$.
@@
(4.24) $\begin{split}2\sum_{i=1}^{k}m_{i}\leq n(r,\infty,U)\leq 2\sum_{i=1}^{k}m_{i}.\end{split}$
@@
Then we estimate the number of zeros of $F$. To this end, we first estimate the number of zeros of $g_{m,n}+1$ for a fixed pair $(m,2n)$ of two nonnegative integers $m$ and $n$. Recall from [9, p. 7] that the Nevanlinna characteristic of the function $\exp(e^{z})$ satisfies
(4.26) $\begin{split}T(r,\exp(e^{z}))\sim\frac{e^{r}}{(2\pi^{3}r)^{-1/2}},\quad r\to\infty.\end{split}$
By definition we see that the coefficients $A_{i}$ and $B_{j}$ in $g_{m,n}$ in (1.5) both satisfy $\sum_{i=0}^{\infty}A_{i}<\infty$ and $\sum_{j=0}^{\infty}|B_{j}|<\infty$ and also that $\sum_{i=0}^{\infty}iA_{i}<\infty$ and $\sum_{j=0}^{\infty}|jB_{j}|<\infty$. Write $g_{m,n}$ in (1.5) as $g_{m,n}(z)=\mathcal{R}(z)\exp(e^{z})$ for a function $\mathcal{R}(z)=R_{m,n}(e^{z})\exp(e^{z})$. Then the first main theorem implies that
@@
as $|z|\to\infty$ outside a set $\Omega_{4}$ of finite linear measure. Then the arguments for determining the order of $A(z)$ are similar as that in the proof of Theorem 1.3. Together with (4.23), the lemma on the logarithmic derivative now implies that $E=F/F^{\prime}$ satisfies
@@
Thus $\rho(A)\leq\rho$. On the other hand, by the relation in (1.8) and the explicit expressions of $\bar{g}_{m_{k},n_{k}}$ in (4.19), (4.20) and (4.21) and also $\tau$ in (4.23) we easily see that $|A(z)|\geq d_{3}|z|^{\rho}$ for some positive constant $d_{3}$ and all $z$ on the positive real axis. Thus we also have $\rho(A)\geq\rho$. Hence $\rho(A)=\rho$.
To complete the proof under the assumption $\rho\in(1/2,1)$, we prove that the second order linear differential equation (1.1) with the constructed coefficient $A(z)$ cannot have any other solution $f_{3}$ linearly independent from $f_{1}$ such that $\lambda(f_{3})<\infty$. Otherwise, there are two nonzero constants $d_{4}$ and $d_{5}$ such that $f_{3}=d_{4}f_{1}+d_{5}f_{2}$ and thus
@@
where $\tau$ is defined in (4.23). By the explicit expressions of $\bar{g}_{m_{k},n_{k}}$ in (4.19), (4.20) and (4.21) we see that the zeros $\check{n}(r,0,f_{3})$ of $f_{3}$ in the lower half-plane satisfies $\check{n}(r,0,f_{3})\geq\exp(d_{6}r^{\rho})$ for some positive constant $d_{6}$ and all large $r$, which implies that $\lambda(f_{3})=\infty$, a contradiction. Thus the desired result follows.
Finally, recalling that $m_{k}=n_{k}=0$ for $k=1,2,3$ in the two sequences $(m_{k})$ and $(n_{k})$, we may use the same arguments as in [5, Section 5] together with the comments in Subsection 4.3 to extend the above results to the case that $\rho\in(n/2,n)$ for any positive integer $n$ by defining two similar functions $G_{0}(z)$ and $G_{1}(z)$ as there, which then yields a locally univalent meromorphic function $F$. The numbers of poles and zeros of $F$ can be estimated similarly as in the case $\rho\in(1/2,1)$ and the order of the corresponding coefficient $A(z)$ of the second order differential equation (1.1) can also be determined. Also, in certain region of the plane we may use similar arguments as for equation (4.38) to show that any other solution $f_{3}$ linearly independent from $f_{1}$ of the second order differential equation (1.1) with the constructed coefficient $A(z)$ cannot satisfy $\lambda(f_{3})<\infty$. We shall not repeat the details. This completes the proof of Theorem 1.6.
@@
$\check{\mathbf{\Pi}}_{k}=\{z=x+iy:\quad y\leq 0,\quad 2\pi\mathcal{N}_{-k}\leq y\leq 2\pi\mathcal{N}_{-k+1}\}$
@@
(5.3) $\begin{split}(\bar{g}_{m_{k},n_{k}}\circ\chi_{m_{k},n_{k}})(z)=\left\{\begin{array}[]{ll}g_{m_{k},n_{k}}\left(\frac{x}{l}+i\frac{-y+2\pi\mathcal{N}_{-k+1}}{N_{k}}\right),&x\geq 0,\\ g_{m_{k},n_{k}}\left(\frac{x+i-y+2i\pi\mathcal{N}_{-k+1}}{N_{k}}\right),&x\leq 0.\end{array}\right.\end{split}$
In the above new definitions, we shall choose $l=1$ in the case (I) and $l=3$ in the case (II) and $l=4$ in the case (III). Note that $n_{k}=0$ for $k=1,2,3$ and thus we have the expressions in (5.2) in $\check{\mathbf{\Pi}}_{1}$, $\check{\mathbf{\Pi}}_{2}$ and $\check{\mathbf{\Pi}}_{3}$. Also note that the two expressions in each of the definition of $\bar{g}_{m_{k},n_{k}}\circ\chi_{m_{k},n_{k}}$ in (5.1), (5.2) or (5.3) coincide on the imaginary axis.
In view of these definitions, the maps $U$, $V$ and $G$ will not commute with complex conjugation, but we shall only define them in the upper half-plane since the definitions in the lower half-plane are similar.
@@
$\begin{split}\mathbf{\Pi}_{k}^{+}=\{x+iy:x>0,\quad 2\pi\mathcal{N}_{k-1}<y<2\pi\mathcal{N}_{k}\}.\end{split}$
@@
(5.5) $\begin{split}G(z)=F(\tau(z))\ \ \text{and}\ \ \tau(z)\sim z\ \ \text{as}\ z\to\infty.\end{split}$
Then $F=G\circ\tau^{-1}$ is a locally univalent meromorphic function. Define $E=F/F^{\prime}$. Then $E$ is a Bank–Laine function. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A(z)$.
@@
Note that $\mathcal{N}_{k}\sim lk$ as $k\to\infty$ by (3.31). We recall that, in the case $10\leq\lambda_{1}\leq\lambda_{2}<$ 1, in the proof of Lemma 3.1 the subsequence $(m_{k_{i}})$ is chosen to satisfy the relation in (3.5) so that the double inequalities in (3.6) or (3.7) hold. For the two sequences $(m_{k})$ and $(n_{k})$ we choose, this implies that $\log\hat{n}(r,0,U_{I})\asymp j=j(r)$ for the greatest $j$ such that $n_{k_{j}}=1$ in the case $\lambda_{2}<1$ and $\log\hat{n}(r,0,U_{I})\asymp r$ in the case $\lambda_{2}=1$. It follows that $\log\hat{n}(r,0,U_{I})\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log\hat{n}(r,0,U_{I})\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Such estimates also hold for $\log\hat{n}(r,0,U_{J})$, $\log\hat{n}(r,0,V_{I})$ and $\log\hat{n}(r,0,V_{J})$. Recall that $2\sum_{j=1}^{k}n_{j}\sim(2\pi)^{-1}(\log(2\pi k))^{2}$ as $k\to\infty$ when $\lambda_{2}=0$ and that $2\sum_{j=1}^{k}n_{j}\sim(2\pi)^{\lambda_{2}-1}k^{\lambda_{2}}$ as $k\to\infty$ when $0<\lambda_{2}\leq 1$, which are pointed out in Subsection 3.4. Then the double inequalities in (5.6) and (5.7) imply that the number of zeros of $G$ in the upper half-plane satisfies $\log\hat{n}(r,0,G)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\hat{n}(r,0,G)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Now the relations in (5.5) imply that the number of zeros of $F$ in the upper half-plane satisfies $\log\hat{n}(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\hat{n}(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Also, for the number of zeros of $F$ in the lower half-plane, we can use the same method as in Subsection 3.4 to show that $\check{n}(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\check{n}(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. By combining the two estimates for $\hat{n}(r,0,F)$ and $\check{n}(r,0,F)$ together, we conclude that $\log n(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log n(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$ and thus $\lambda(f_{2})=\infty$.
Moreover, by similar arguments as in Subsection 3.4 we can estimate the modulus of $\log\log|G|$ and $\log\log|F|$ to conclude that the resulting Bank–Laine function $E$ satisfies $m(r,1/E)=O(r)$ and further that $\rho(A)\leq 1$. On the other hand, by considering the modulus of $E$ on the positive real axis and the negative real axis respectively we may conclude that $\rho(A)\geq 1$. Hence $\rho(A)=1$. Also, for any other solution $f_{3}$ linearly independent from $f_{1}$ of the second order differential equation (1.1) with the constructed coefficient $A(z)$ we may consider the equation (4.38) and show that $f_{3}$ cannot satisfy $\lambda(f_{3})<\infty$. This completes the proof for the case $n=1$.

### v3

@@ 4.2. Definition of a quasiregular map
of the upper half-plane we will consider the functions
(4.19) $\bar{g}_{m_{k},n_{k}}(z)=g_{m_{k},n_{k}}\left(x+i\frac{y-2\pi N_{k-1}}{k}\right)$
@@
(4.20) $\bar{g}_{m_{k},n_{k}}(z)=\frac{1}{2}\left[g_{m_{k},n_{k}}\left(x+i\frac{y-2\pi N_{k-1}}{k}\right)+1\right]$
@@
$\check{\Pi}_{-k}=\{z=x+iy:\quad 2\pi(-k)\leq y\leq 2\pi(-k+1)\}$
of the lower half-plane we will consider the functions
(4.21) $\bar{g}_{m_{k},n_{k}}(z)=g_{m_{k},n_{k}}\left(x-i\frac{y-2\pi N_{1-k}}{k}\right).$
In view of the explicit expressions of $\bar{g}_{m_{k},n_{k}}$ defined in (4.19), (4.20) and (4.21), the maps $U$, $V$ and $G$ will not commute with the complex conjugation. However, we shall only define them in the upper half-plane since the definitions in the lower half-plane are similar.
@@
$\begin{split}\hat{\Pi}_{k}^{+}=\{x+iy:x>0,\quad 2\pi(k-1)<y<2\pi k\}.\end{split}$
@@ 4.3. Completion of the proof
With the definition of $G$ in (4.22) in previous section, here we first prove Theorem 1.6 under the assumption that $\rho\in(1/2,1)$. Now, we can follow exactly the same process as that in [5, Section 3.3] to estimate the dilation of $G$ and thus prove that $G$ is indeed quasiregular in the plane and also that $G$ satisfies the hypothesis of the Teichmüller–Wittich–Belinskii theorem. This theorem, together with the existence theorem for quasiconformal mappings, yields that there exists a quasiconformal homeomorphism $\tau:\mathbb{C}\to\mathbb{C}$ and a meromorphic function $F$ such that
@@
Then $F=G\circ\tau^{-1}$ is a locally univalent meromorphic function. Define $E=F/F^{\prime}$. Then $E$ is a Bank–Laine function. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A$.
Then we use similar arguments as in the proof of Theorem 1.3 to estimate the numbers of poles and zeros of $F$ respectively and also to determine the orders of $E$ and $A$. Again, two positive quantities $S_{1}(r)$ and $S_{2}(r)$ are written as $S_{1}(r)\asymp S_{2}(r)$ if there are two positive constants $d_{1}$ and $d_{2}$ such that $d_{2}S_{2}(r)\leq S_{1}(r)\leq d_{1}S_{2}(r)$ for all large $r$.
@@
(4.24) $\begin{split}\sum_{i=1}^{k}m_{i}\leq n(r,\infty,U)\leq 2\sum_{i=1}^{k}m_{i}.\end{split}$
@@
Then we estimate the number of zeros of $F$. To this end, we first estimate the number of zeros of $g_{m,n}+1$ for a fixed pair $(m,2n)$ of two nonnegative integers $m$ and $n$. Recall from [10, p. 7] that the Nevanlinna characteristic of the function $\exp(e^{z})$ satisfies
(4.26) $\begin{split}T(r,\exp(e^{z}))\sim\frac{e^{r}}{(2\pi^{3}r)^{1/2}},\quad r\to\infty.\end{split}$
By definition we see that the coefficients $A_{i}$ and $B_{j}$ in $g_{m,n}$ in (1.5) both satisfy $\sum_{i=0}^{\infty}A_{i}<\infty$ and $\sum_{j=0}^{\infty}|B_{j}|<\infty$ and also that $\sum_{i=0}^{\infty}iA_{i}<\infty$ and $\sum_{j=0}^{\infty}|jB_{j}|<\infty$. Write $g_{m,n}$ in (1.5) as $g_{m,n}(z)=\mathcal{R}(z)\exp(e^{z})$ for a function $\mathcal{R}(z)=R_{m,n}(e^{z})$. Then the first main theorem implies that
@@
as $|z|\to\infty$ outside a set $\Omega_{4}$ of finite linear measure. Together with (4.23), the lemma on the logarithmic derivative now implies that $E=F/F^{\prime}$ satisfies
@@
Thus $\rho(A)\leq\rho$. On the other hand, we note that $g_{1,1}$ is analytic in $\hat{\Pi}_{1}^{+}$ as well as in $\check{\Pi}_{-1}^{+}$. Let $\bar{\eta}:H^{+}\to H^{+}$ be the map such that $\bar{\eta}=z^{\rho}$. Denote $\bar{Q}=Q+\bar{s}_{1,1}$. Note that $\bar{\eta}$ and $\bar{Q}$ both commute with the complex conjugation and also that are both analytic for $z$ in the right half-plane such that $\text{Re}\ z\geq 1$. Then, for $\tau$ in (4.23), by the same arguments as in the proof of Theorem 1.3 we may choose a curve in $(\tau\circ\bar{\eta}^{-1}\circ\bar{Q}^{-1})(\hat{\Pi}_{1}^{+}\cup\check{\Pi}_{-1}^{+})$, say $\mathcal{L}^{+}$, such that $(\bar{Q}\circ\bar{\eta}\circ\tau^{-1})(\mathcal{L}^{+})$ is the positive real axis and also that $\tau$ satisfies $|\tau^{\prime}(z)|\leq 2|z|$ for all $z\in(\bar{\eta}^{-1}\circ\bar{Q}^{-1})(\mathcal{L}^{+})$ such that $|z|$ is large. Then, by the relation in (1.8) as well as the definitions of $u_{m,n}$ in Subsection 4.2 and the relation $F=G\circ\tau^{-1}$, we easily see that $|A|\geq d_{3}|z|^{\rho}$ for some positive constant $d_{3}$ and all $z\in\mathcal{L}^{+}$ such that $|z|$ is large, implying that $\rho(A)\geq\rho$. Hence $\rho(A)=\rho$.
To complete the proof under the assumption $\rho\in(1/2,1)$, we show that the second order differential equation (1.1) with the constructed coefficient $A$ cannot have any other solution $f_{3}$ linearly independent from $f_{1}$ such that $\lambda(f_{3})<\infty$. Otherwise, there are two nonzero constants $d_{4}$ and $d_{5}$ such that $f_{3}=d_{4}f_{1}+d_{5}f_{2}$ and thus
@@
where $\tau$ is defined in (4.23). By the explicit expressions of $\bar{g}_{m_{k},n_{k}}$ in (4.19), (4.20) and (4.21) we see that the zeros $\check{n}(r,0,f_{3})$ of $f_{3}$ in the lower half-plane satisfies $\log\check{n}(r,0,f_{3})\geq d_{6}r^{\rho}$ for some positive constant $d_{6}$ and all large $r$, which implies that $\lambda(f_{3})=\infty$, a contradiction. Thus the desired result follows.
Finally, recalling that $m_{k}=n_{k}=0$ for $k=1,2,3$ in the two sequences $(m_{k})$ and $(n_{k})$, we may follow exactly the same process as that in [5, Section 5] together with the comments in Subsection 3.5 to extend the above results to the case $\rho\in(n/2,n)$ for any positive integer $n$, except that the corresponding two functions $G_{0}$ and $G_{1}$ there are now defined by using the function $G$ in the case $\rho\in(1/2,1)$ here. Then we obtain a locally univalent meromorphic function $F$. The numbers of poles and zeros of $F$ can be estimated in a similar way as in the case $\rho\in(1/2,1)$ and the order of the corresponding coefficient $A$ of the second order differential equation (1.1) can also be determined by using similar arguments as there. Moreover, in certain region of the plane we may use similar arguments as for equation (4.38) to show that any other solution $f_{3}$ linearly independent from $f_{1}$ of the second order differential equation (1.1) with the constructed coefficient $A$ cannot satisfy $\lambda(f_{3})<\infty$. We shall not repeat the details. This completes the proof of Theorem 1.6.
@@
    if $\lambda_{1}\leq\lambda_{2}<1$, then we choose the sequence $(m_{k})$ and $(n_{k})$ constructed in Subsection 3.1 respectively; moreover, by slightly modifying the proof of Lemma 3.1, we may choose a different sequence $(n_{k})$ from the one there so that the function $\tilde{h}_{2}$ related to $\lambda_{2}$ can also satisfy $\tilde{h}_{2}(x)=(1+o(1))\lambda_{3}\log r$ for any given number $\lambda_{3}\in[0,\infty)$ in the case $\lambda_{2}=0$;
@@
$\check{\mathbf{\Pi}}_{-k}=\{z=x+iy:\quad y\leq 0,\quad 2\pi\mathcal{N}_{-k}\leq y\leq 2\pi\mathcal{N}_{-k+1}\}$
@@
(5.3) $\begin{split}(\bar{g}_{m_{k},n_{k}}\circ\chi_{m_{k},n_{k}})(z)=\left\{\begin{array}[]{ll}g_{m_{k},n_{k}}\left(\frac{x}{l}-i\frac{y-2\pi\mathcal{N}_{1-k}}{N_{k}}\right),&x\geq 0,\\ g_{m_{k},n_{k}}\left(\frac{x-iy+2i\pi\mathcal{N}_{1-k}}{N_{k}}\right),&x\leq 0.\end{array}\right.\end{split}$
In the above new definitions, we shall choose $l=1$ in the case (I) and $l=3$ in the case (II) and $l=4$ in the case (III). Note that $n_{k}=0$ for $k=1,2,3$ and thus we have the expressions in (5.2) in $\hat{\mathbf{\Pi}}_{1}$, $\hat{\mathbf{\Pi}}_{2}$ and $\hat{\mathbf{\Pi}}_{3}$. Also note that the two expressions in each of the definition of $\bar{g}_{m_{k},n_{k}}\circ\chi_{m_{k},n_{k}}$ in (5.1), (5.2) or (5.3) coincide on the imaginary axis.
In view of these definitions, the maps $U$, $V$ and $G$ will not commute with the complex conjugation, but we shall only define them in the upper half-plane since the definitions in the lower half-plane are similar.
@@
$\begin{split}\hat{\mathbf{\Pi}}_{k}^{+}=\{x+iy:x>0,\quad 2\pi\mathcal{N}_{k-1}<y<2\pi\mathcal{N}_{k}\}.\end{split}$
@@
Then $F=G\circ\tau^{-1}$ is a locally univalent meromorphic function. Define $E=F/F^{\prime}$. Then $E$ is a Bank–Laine function. Clearly, $E=f_{1}f_{2}$ for two normalized solutions of the second order differential equation (1.1) with an entire coefficient $A$.
@@
Note that $\mathcal{N}_{k}\sim lk$ as $k\to\infty$ by (3.31). We recall that, in the case $10\leq\lambda_{1}\leq\lambda_{2}<$ 1, in the proof of Lemma 3.1 the subsequence $(m_{k_{i}})$ is chosen to satisfy the relation in (3.5) so that the double inequalities in (3.6) or (3.7) hold. For the two sequences $(m_{k})$ and $(n_{k})$ we choose, this implies that $\log\hat{n}(r,0,U_{I})\asymp j=j(r)$ for the greatest $j$ such that $n_{k_{j}}=1$ in the case $\lambda_{2}<1$ and $\log\hat{n}(r,0,U_{I})\asymp r$ in the case $\lambda_{2}=1$. It follows that $\hat{n}(r,0,U_{I})\asymp r^{\lambda_{3}}$ or $\log\hat{n}(r,0,U_{I})\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log\hat{n}(r,0,U_{I})\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Such estimates also hold for $\log\hat{n}(r,0,U_{J})$, $\log\hat{n}(r,0,V_{I})$ and $\log\hat{n}(r,0,V_{J})$. Recall that $2\sum_{j=1}^{k}n_{j}\sim(2\pi)^{-1}(\log(2\pi k))^{2}$ as $k\to\infty$ when $\lambda_{2}=0$ and that $2\sum_{j=1}^{k}n_{j}\sim(2\pi)^{\lambda_{2}-1}k^{\lambda_{2}}$ as $k\to\infty$ when $0<\lambda_{2}\leq 1$, which are pointed out in Subsection 3.4. Then the double inequalities in (5.6) and (5.7) imply that the number of zeros of $G$ in the upper half-plane satisfies $\hat{n}(r,0,G)\asymp r^{\lambda_{3}}$ or $\log\hat{n}(r,0,G)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log\hat{n}(r,0,G)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Now the relations in (5.5) imply that the number of zeros of $F$ in the upper half-plane satisfies $\hat{n}(r,0,G)\asymp r^{\lambda_{3}}$ or $\log\hat{n}(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log\hat{n}(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. Also, for the number of zeros of $F$ in the lower half-plane, we can use the same method as in Subsection 3.4 to show that $\check{n}(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\check{n}(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$. By combining the two estimates for $\hat{n}(r,0,F)$ and $\check{n}(r,0,F)$ together, we conclude that $n(r,0,G)\asymp r^{\lambda_{3}}$ or $\log n(r,0,F)\asymp(\log r)^{2}$ when $\lambda_{2}=0$ and $\log n(r,0,F)\asymp r^{\lambda_{2}}$ when $0<\lambda_{2}\leq 1$ and thus $\lambda(f_{2})=\infty$.
Moreover, by similar arguments as in Subsection 3.4 we can estimate the modulus of $\log\log|G|$ and $\log\log|F|$ to conclude that the resulting Bank–Laine function $E$ satisfies $m(r,1/E)=O(r)$ and further that $\rho(A)\leq 1$. On the other hand, by considering the modulus of $E$ on the positive real axis we may conclude that $\rho(A)\geq 1$. Hence $\rho(A)=1$. Also, for any other solution $f_{3}$ linearly independent from $f_{1}$ of the second order differential equation (1.1) with the constructed coefficient $A$ we may consider the equation (4.38) and show that $f_{3}$ cannot satisfy $\lambda(f_{3})<\infty$. This completes the proof for the case $n=1$.